OpenStudy (pokemon23):

Hello openstudy I just got back my exam and I got a 99 but their a question I got wrong that cost me some points the question is Josh is 2 years older than Paul. The sum of their ages is 24. If p represents Paul age.

OpenStudy (pokemon23):

represent Josh's age in terms of p my answer is 2+p

OpenStudy (pokemon23):

this question consist of three parts

OpenStudy (pokemon23):

write an equation thaT SAYS THE SUM of their ages is 24

OpenStudy (pokemon23):

solve your equation .

OpenStudy (pokemon23):

old is each is each boy ? Josh:_____ and Paul _____ years old

OpenStudy (pokemon23):

my equation was p+2=24

OpenStudy (s):

Paul = p years old Josh = p + 2 years old (p) + (p + 2) = 24 2p + 2 = 24 2p = 22 p = 11 So Paul is 11 years old And Josh is 11 + 2 = 13 years old

OpenStudy (pokemon23):

thanks